OpenStudy (anonymous):

what are the possible dimensions for a rectangle when the area is 8x^2-3x-8x+3

OpenStudy (anonymous):

Factor the trinomial; since a + b + c = 0 then the quadratic equation f(x) = 0 has 2 real roots x = 1 and x = c/a = 3/8. Factored form: y = (x - 1)(8x - 3)

OpenStudy (anonymous):

Ayyee, thanks bro